OpenStudy (evonhowell):

6 (3-x) = -4 (2x-1) Find x Value The X Value is 7 or -7 just can't figure out how to get the answer

OpenStudy (evonhowell):

And I got alot more of this crap and my teacher didn't tell me how to do them

OpenStudy (anonymous):

6(3- x) = -4(2x - 1) the first thing you want to do is distribute the 6 and the -4 through the parenthesis. Can you do that ?

OpenStudy (evonhowell):

I know what distribute means but don't know how to do it exactly

OpenStudy (anonymous):

6(3 - x) = 6 * 3 - 6 * x and -4(2x - 1) = -4 * 2x - (-4) * 1 understand ?

OpenStudy (anonymous):

let me restate that second one....-4 * 2x + -4 * -1

OpenStudy (evonhowell):

soo 12x = 13x?

OpenStudy (anonymous):

6(3 - x) = -4(2x - 1) 18 - 6x = -8x + 4 can you take it from here because I have to go...sorry

OpenStudy (anonymous):

18 - 6x = -8x + 4 subtract 18 from both sides -6x = -8x + 4 - 18 -6x = -8x - 14 now add 8x to both sides -6x + 8x = -14 2x = -14 x = -7 sorry I couldn't explain better, I have to go now

OpenStudy (evonhowell):

-6x - 90 = 9x

OpenStudy (evonhowell):

Tho I am so lost with all this stuff

thomaster (thomaster):

subtract 9x on both sides and add 90 on both sides You'd get -15x=90 now divide by -15

thomaster (thomaster):

The point is that you bring the variables (x in this case) to the left side of the equation, and the numbers to the right side.

OpenStudy (evonhowell):

-6 is the answer?

thomaster (thomaster):

\(\large\dfrac{~~90}{-15}=-6\) so yea

OpenStudy (evonhowell):

so whats about 3x - 4 = 10-4x How would i do this one

thomaster (thomaster):

bring numbers to the right, x to the left. To bring x to the left, you would have to add 4x on both side. That way, you can cancel the -4x on the right. \(\large3x - 4 = 10-4x\) \(\large3x~\underline{+4x} - 4 = 10-4x~\underline{+4x}\) \(\large7x - 4 = 10\)

thomaster (thomaster):

Now you only have to bring -4 to the right Just add 4 on both sides and you'll cancel the -4 on the left. So it'll become: 7x=14 now you divide 14 by 7 to solve for x

OpenStudy (evonhowell):

x=2

thomaster (thomaster):

exactly :)

OpenStudy (evonhowell):

can we do 2 more?

thomaster (thomaster):

sure

OpenStudy (evonhowell):

-7 - (-4x) = -4x+9

OpenStudy (evonhowell):

oops 2 more after this 1 my bad

thomaster (thomaster):

Alright first of all: -(-4x) is the same as +4x

thomaster (thomaster):

So the equation would be -7 +4x = -4x+9

thomaster (thomaster):

now we want that -7 on the other side of the equation, so we add 7 on both sides 4x=-4x+16

thomaster (thomaster):

last thing is to add 4x on both sides to get that -4x to the left 8x=16 Do you know what to do now?

OpenStudy (evonhowell):

Divide both sides by 8?

thomaster (thomaster):

Yep :)

OpenStudy (evonhowell):

x=2

thomaster (thomaster):

correct

OpenStudy (evonhowell):